A continuous random variable has a density functionš‘“(š‘„) = 2(5 ā€“ x) 5 , where 2<x<3. Calculate the following probability correct up to 3 decimal places, and make the graph for part (a) only in Answer sheet:

P (x < 2.5)

P (x > 2.2)

P (2.1 ā‰¤ š‘„ ā‰¤ 2.7)


1
Expert's answer
2020-11-30T13:34:45-0500

I) p(x<2.5)= "\displaystyle\int\limits^{{2.5}}_{{2}} 2\cdot\dfrac{(5-x)}{5}\,{{d}x}"


= "\int_{2}^{2.5}2dx-\int_{2}^{2.5}\frac{2x}{5}dx"


"=2x-\frac{x^2}{5}| 2 to 2.5"

"=\frac{11}{20}"


ii)"\displaystyle\int\limits^{{3}}_{{2.2}} 2\cdot\dfrac{(5-x)}{5}\,{{d}x}"


"\int_{2.2}^{3}2dx-\int_{2.2}^{3}\frac{2x}{5}dx"


"=\frac{96}{125}"


iii)"\displaystyle\int\limits^{{2.7}}_{{2.1}} 2\cdot\dfrac{(5-x)}{5}\,{{d}x}"


"\int_{2.1}^{2.7}2dx-\int_{2.1}^{2.7}\frac{2x}{5}dx"


"=\dfrac{78}{125}"